OpenStudy (anonymous):

-170 = 500/1-4e^-2k HELP SOLVE PLEASE!

OpenStudy (anonymous):

k = (335 e^2)/2

OpenStudy (anonymous):

?how did you get 335

OpenStudy (anonymous):

could you repeat the equation again but with parenthesis?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-170 = (500)/(1-4e^-2k)

OpenStudy (anonymous):

k = (67 e^2)/68

OpenStudy (anonymous):

-170(1 - 4e^-2k) = 500 thus 1 - 4e^-2k = -500/170 thus 1 - 4e^-2k = -50/17 thus 4e^-2k = 1 + 50/17 thus e^-2k = (1 + 50/17)/4 thus e^-2k = (67/68) thus -2k = ln(67/68) thus k = -1/2 * ln(67/68)

OpenStudy (anonymous):

ok, thank you!